# The Junction of Immersive Analytics and Virtual Reconstructions – A Case Study on the Mausoleum of Emperor Maxentius thanks: DFG, project ID 251654672 – TRR 161

### v2

@@ Abstract
Virtual Archaeology has significantly evolved over the last decades through advancements in data acquisition and representation, for example, by improved data recording technologies and virtual reality devices. Immersive environments provide novel ways to present historical events or objects with high visual quality for both the general public and researchers. Here, we examine how the emerging field of Immersive Analytics can contribute to enhancing the understanding and exploration of archaeological data and explore the junction of Virtual Archaeology and Immersive Analytics. We discuss a selection of features already used by the community and look into how optimizing those can facilitate the discourse and impression of cultural heritage objects. As an argumentational basis, we introduce and utilize three digital reconstruction interpretations of the mausoleum of the late Roman emperor Maxentius in Rome, which are based on prior scientific work and a typological embedding. Based on our work, we advocate the value of combining historical and computer science expertise to optimize immersive environments for virtual reconstructions to facilitate deeper understanding and interactive exploration of archaeological data.
Index Terms:
virtual archaeology; cultural heritage; 3D reconstruction; immersive analytics; virtual reality; Maxentius mausoleum;
@@ I Introduction
Terms like Virtual Archaeology, Virtual Restoration, and Virtual Reconstruction date back to the 1990s, when researchers investigated the potential of virtualization for work on cultural heritage [1, 2]. Throughout the years, Virtual Archaeology has since grown into an entire field of research and seen strong development in its technological capabilities in terms of data acquisition and data representation, as well as definitions of charters that employ meaningful standards for visualizations and virtual reconstructions [3, 4]. According to the Seville Charter [3], Virtual Archaeology concerns itself with finding comprehensive ways of visualizing cultural heritage objects.
@@
The visualization of archaeological objects began in the 2D realm and has since increasingly seen an extension through 3D to stereoscopic 3D (s3D) applications (e. g., models in virtual reality) since 2D visualizations often fall short of capturing the complexity and spatial relationships inherent in archaeological data. Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino [5]. Most commonly, immersive technologies are understood to be either virtual reality (VR) as an entirely virtual environment, augmented reality (AR) as an overlay of additional information into the real world, or mixed reality (MR) as a blend between the virtual and real world. By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artifacts but also reconstructed objects, background information, or a contextually appropriate environment.
@@
Rome, as a focal point of archaeological work in Europe, enjoys high popularity for such work, leading to an increase in more extensive reconstruction visualizations. An example of such an immersive model representation is Rome Reborn, a virtual reality project that has been running since 1996 at the University of Indiana, which reconstructs Rome around 320 AD [6, pp. 23–25]. The transparency of the underlying archaeological data origin of such public audience-oriented applications tends to be selective, and their authenticity and reliability are therefore limited. This is only natural because it can not be the goal of an immersive environment targeted at the broader public to discuss different opinions of scholarship regarding the reconstruction [7, pp. 13–14].
@@
To support scientific debates, digital reconstructions are useful for developing innovative questions on the historical perception and sociopolitical embedding of monuments and objects, visualizing different chronological phases of archaeological sites, and reflecting on historical monuments’ construction history and building process. A clear indication of different phases, information sources, or certainty of validity can, among others, take the form of color mappings or opacity of the respective components integrated into the visualisation [8, 9, 10].
@@
To combine the potential of immersive environments and scientific debates, the relatively young field of Immersive Analytics (IA) offers plenty of advancements. IA uses engaging and embodied tools to support data understanding [11], explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored towards the use in specific application areas. IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality. As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it [12, 13]. The investigation of which navigation or interaction methods are most appropriate for a specific context [14], how effective VR is in teaching contexts [15], or the impact of different modalities like audio or haptics [16, 17] are some of its core questions.
@@
In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, such as by investigating how to annotate specific objects or positions or how to collaborate in which contexts. Creating immersive environments that allow people to explore sites, artifacts, and reconstructed objects virtually, to add annotations and landmarks for idea communication, and to obtain more than mere visual impressions of cultural heritage objects supports bridging the gap between academia and the public. Thus, it is crucial to emphasize the optimization of those supporting features based on the cultural heritage item displayed and the user group towards which the virtual environment is targeted.
@@
Since immersive experiences are broadly utilized for dissemination, like museums or teaching, as well as for research [18, 19], and evidence suggests that discourse in Virtual Archaeology is mainly driven by advancing technologies [20], we discuss perspectives of how the mentioned potentials can be utilized. We discuss how Immersive Analytics methods can be used for Virtual Archaeology, specifically for virtual reconstructions. We use three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig. 1). They are based on ongoing scientific debates about what the mausoleum might have looked like in its prime. We provide these reconstruction models in addition to software to inspect them in VR.
@@
Thus, our contribution is comprised of a perspective on how Immersive Analytics research can help to assist archaeological discourse and the three digital reconstructions of the Maxentius mausoleum that we created. Furthermore, we provide an example software to demo some of the possible concepts and techniques.
@@
Figure 1: This image shows the three proposed models of the Mausoleum of Maxentius we discuss in this work. The first model (Maxentius A) closely follows Rasch’s reconstruction proposal [21]. The second model (Maxentius B) is based on Rasch’s model but has a closed dome and seven windows in the clerestory modelled after the Tor de’ Schiavi [22]. The third model (Maxentius C) tries to merge the dome and the windows of Tor de’ Schiavi to the proportions of the Maxentius Mausoleum.
@@ II Background
The popularity of immersive technologies, such as AR and VR, for museums or expositions, teaching, and even research in archaeological contexts has grown [18, 19, 6, 23, 24]. As we focus on reconstructions in this work and its interdisciplinary nature, it’s important to clarify the different types of reconstructions in the digital realm. Digital reconstruction involves creating objects using computational tools, while virtual reconstruction is the use of virtual models to recover a building or man-made object visually [3]. Immersive reconstruction extends this further by embedding objects in environments along Milgram’s reality-virtuality continuum [5], such as AR or VR, to enhance user experience through interaction and context.
As is the nature of a reconstruction, it is not granted that all pieces of a monument, for instance, are known. Researchers have to complement the known parts with modeled ones, whose form and structure derive from different sources, like paintings, folklore, derivation from existing pieces, and so on. These different sources need to be pointed out, as Demetrescu describes in his work [9]. The sources of the reconstruction components, as well as the surrounding area and information on the object’s purpose or its material, are all data that can be transported via various means, especially so with the help of immersive technologies.
Immersive technologies can support reasoning [25] by embedding data into the physical world [26], the recall of data [27], or spatial judgment accuracy [28]. This is also relevant in Virtual Archaeology, as such environments have proven to create a heightened sense of presence in its’ context, as already shown by Kyrlitsias, Christofi, Michael-Grigoriou, Banakou, and Ioannou [29]. Presence, as a form of psychological immersion, can also be called engagement [30], and is an important concept, as there is evidence that task performance and recall are boosted through heightened engagement [31].
The research direction that concerns itself with investigating how to exploit those engaging tools to increase data understanding is called Immersive Analytics [11]. In their book on Immersive Analytics, Marriott et al. describe a number of relevant aspects apart from those already discussed. One of those is Situated analytics, which puts data representations in relation to germane objects, places, and persons to improve understanding and decision-making [32]. For example, it might overlay an existing object with analysis information about it and is, thus, tightly linked to AR applications. Another is the Embodiment, which relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance rather than point-and-click interaction as it is typical for desktop PCs [30]. That said, the concept of multi-modality is closely related, as it encompasses input and output addressing multiple senses, such as vision, hearing, touch, kinesthesis, and (less commonly) smell.
@@ III Maxentius Mausoleum
Figure 2: The current state of the Maxentius Mausoleum. Only the Krypta still remains. The front porch is partly missing and has partly been used to build a farmhouse.
@@
The monumental mausoleum was strongly connected to Maxentius’ rule and likely almost completed when Maxentius was killed in 312 AD [22, pp. 92–93]. Its few remaining traces provide an excellent case study for the application of virtual archaeology. Typologically, the mausoleum can be attributed to a type of podium round building, which – according to the architectural historian Jürgen Rasch – ultimately derives from the Pantheon. From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a drawn reconstruction of the building modeled on the Pantheon with its specific architectural components: A podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey [34, p. 84]. However, this Pantheon version of the Maxentius Mausoleum has been questioned by Mark Johnson, referring to a temporally and typologically similar building – another late antique imperial mausoleum at the via Praenestina (the so-called Mausoleum of Tor de’ Schiavi) [34]. This monument is located, like the Maxentius Mausoleum, on the outskirts of Rome, close to an imperial villa suburbana. It follows the design of a podium round building featuring a slightly smaller, tetrastyle front porch with a domed vault without an oculus but four windows in the clerestory. These different perspectives, into which Maxentius’ Mausoleum can be classified, result in different conceivable models for its reconstruction, which will be discussed below.
@@ IV Immersive Analytics for the Maxentius Mausoleum
Regarding the Maxentius Mausoleum, we chose Rasch’s work as a foundation for a first reconstruction (Maxentius A), which closely follows his suggested measurements. The second model (Maxentius B) assumes a closed dome following Johnson’s proposal [22, pp. 91-92] and adds seven round windows in the clerestory with a diameter of 1.80 meters considering the proportions of Tor de’ Schiavi [34, p. 28]. The third model (Maxentius C) adapts the flatter dome of Tor de’ Schiavi [34, p. 68] and the windows with a larger diameter. This model visualizes, as suggested by Johnson, a version of the Maxentius Mausoleum modeled closely on Tor de’ Schiavi [22, p. 92]. Consequently, model B can be considered a merger of models A and C. Thus, we have three reconstruction models of the same object (see Fig. 1), which mainly differ in the dome’s design and allow us to discuss which of the hypotheses about the part of the mausoleum is most plausible. We created the reconstructions with the 3D modeling software Blender. The components of the models consist of combined simple geometric forms, which were manipulated through manual node shifting to obtain necessary details. For each sub-component, once constructed from primitives, the measurements for those components were linearly scaled and applied. Thanks to the precise measurements we get from Rasch’s work, the models provide correct proportions, which helps us to obtain a fitting impression of the mausoleum’s former glory. In Fig. 3, we demonstrate this via overlaying Maxentius A transparently over a photogrammetry-based model of the mausoleum acquired from airborne imaging using a drone.
Figure 3: Cross-faded version of Maxentius A. Therefore, the different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
We provide software (see Fig. 4) that presents the three hypothesized models of the Maxentius mausoleum. The software provides a VR environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index. We chose VR as it represents an s3D environment with head-tracking, which allows for improved spatial judgment. Inside the immersive environment, the user will start in front of the Maxentius A stairway. The user can move around, using the continuous movement metaphor, and can snap-turn or move around freely, depending on the physical space available around them.
@@
There is a small panel with three buttons to switch between the three reconstruction models and a toggle to turn on or off a coloring akin to the extended matrix by Demetrescu by touching those buttons with their hands. Switching between models is especially relevant in a research context for comparing several different hypotheses directly while still in the iVR environment. Consequently, in dialogue with domain experts, compiling ways to compare the reconstruction models intuitively or conveniently helps to improve cognitive affordance.
To the right of the panel is a miniature of the model. By holding their virtual hands inside the miniature and performing a grabbing motion, a vibration of the controller indicates interaction, and they can control the light’s angle of incidence on the model. There is a small yellow sphere indicating the relative position of the light source at the miniature. This small sphere is intended to serve as a visualization to ease the intuition of where the light comes from.
This idiom is supposed to help when analyzing how and where sunlight comes from for the mausoleum. For our mausoleum scenario, it would make sense to look into relevant ways to interact with light sources. An optimized light interaction system could lead to a more fluent experience, which in turn may help experts justify the different shapes of the dome and the light conditions inside. Furthermore, users can place small squares to leave small randomly colored markers at points of interest behind. In general, different types of markers could be developed. For example, users could be enabled to colorize specific areas of the wall to mark the locations of the numerous drawings on the walls, which were added long after the mausoleum was originally built. The type of marker, color coding, or maybe even an acoustic signal could then again be investigated for its effectiveness in finding the different spots.
Lastly, the software can also be used without a VR headset with very simple keyboard and mouse controls, such that it can be viewed without the need for additional hardware. We provide the software via a Zenodo upload at [35].

### v3

@@ Abstract
Virtual archaeology has significantly evolved over the last few decades through advancements in data acquisition and representation; for example, by improved data recording technologies and virtual reality devices. Immersive environments provide novel ways to present historical events or objects with high visual quality for both the general public and researchers. Here, we examine how the emerging field of immersive analytics can contribute to enhancing the understanding and exploration of archaeological data, and we explore the junction of virtual archaeology and immersive analytics. We discuss a selection of features already used by the community and examine how optimizing these can facilitate the discourse on cultural heritage objects. As a basis for discussion, we introduce and utilize three digital reconstruction interpretations of the mausoleum of the late Roman Emperor Maxentius in Rome, which are based on prior scientific work and a typological framework. Based on our work, we advocate for the value of combining historical and computer science expertise to optimize immersive environments for virtual reconstructions, thereby facilitating a deeper understanding and interactive exploration of archaeological data.
@@ I Introduction
Terms such as virtual archaeology, virtual restoration, and virtual reconstruction date back to the 1990s, when researchers investigated the potential of virtualization for work on cultural heritage [1, 2]. Over the years, virtual archaeology has evolved into an entire field of research, with significant advancements in its technological capabilities in terms of data acquisition and data representation, as well as the development of charters that employ meaningful principles for visualizations and virtual reconstructions [3, 4]. According to the Seville Charter [3], virtual archaeology concerns itself with finding comprehensive ways of visualizing cultural heritage objects.
@@
The visualization of archaeological objects began in the 2D realm and has since increasingly seen an extension through 3D to stereoscopic 3D (s3D) applications (e. g., models in virtual reality) since 2D visualizations often fall short of capturing the complexity and spatial relationships inherent in archaeological data. Immersive technologies can typically be classified by the reality-virtuality continuum by Milgram and Kishino [5]. Most commonly, immersive technologies are understood to be either virtual reality (VR), which is an entirely virtual environment; augmented reality (AR), which is an overlay of additional information into the real world; or mixed reality (MR), which is a blend between the virtual and real worlds. By leveraging immersive technologies, researchers can generate detailed and realistic 3D experiences of not only the sites and artifacts but also reconstructed objects, background information, or a contextually appropriate environment.
@@
Rome, as a focal point of archaeological work in Europe, enjoys high popularity for such work, leading to an increase in more extensive reconstruction visualizations. An example of such an immersive model representation is Rome Reborn, a virtual reality project that has been running since 1996 at the University of Indiana, which reconstructs Rome around 320 AD [6, pp. 23–25]. The transparency of the underlying archaeological data origin of such public audience-oriented applications tends to be selective, and their authenticity and reliability are therefore limited. This is only natural because the goal of an immersive environment targeted at the broader public is not to discuss different scholarly opinions regarding the reconstruction [7, pp. 13–14].
@@
To support scientific debates, digital reconstructions are useful for developing innovative questions about the historical perception and sociopolitical context of monuments and objects, visualizing different chronological phases of archaeological sites, and reflecting on the construction history and building process of historical monuments. A clear indication of different phases, information sources, or certainty of validity can, among others, take the form of color mappings or opacity of the respective components integrated into the visualization [8, 9, 10].
@@
To combine the potential of immersive environments with scientific debates, the relatively young field of immersive analytics (IA) offers numerous advancements. IA utilizes engaging and embodied tools to support data understanding [11], explores the potential of immersive environments for data analysis, and develops and evaluates designs of such environments tailored for specific application areas. IA combines topics from visual analytics, computer graphics, human-computer interaction, and virtual reality. As a consequence, several areas that exhibit potential for archaeological research, including s3D, collaborative analysis, and multimodal representation and interaction, can be considered a part of it [12, 13]. The investigation of which navigation or interaction methods are most appropriate for a specific context [14], how effective VR is in teaching contexts [15], or the impact of different modalities such as audio or haptics [16, 17] are only some of its research topics.
@@
In addition to enhancing data representation and analysis, IA can help communicate results in a more engaging and accessible way, such as by investigating how to annotate specific objects or positions or how to collaborate in which contexts. Creating immersive environments that allow people to virtually explore sites, artifacts, and reconstructed objects, add annotations and landmarks for idea communication, and obtain more than mere visual impressions of cultural heritage objects supports bridging the gap between academia and the public. Thus, it is crucial to emphasize the optimization of those supporting features based on the cultural heritage item displayed and the user group towards which the virtual environment is targeted.
@@
Since immersive experiences are widely utilized for dissemination, such as in museums or teaching, as well as for research [18, 19], and evidence suggests that discourse in virtual archaeology is primarily driven by advancing technologies [20], we discuss perspectives on how the mentioned potentials can be utilized. We consider how immersive analytics methods can be used for virtual archaeology, specifically for virtual reconstructions. We use three reconstructions of the Maxentius Mausoleum in Rome for our argumentation (see Fig. 1). They are based on ongoing scientific debates about what the mausoleum might have looked like in its prime. We provide these reconstruction models, along with software, to inspect them in VR.
@@
Thus, our contribution consists of a perspective on how immersive analytics research can assist archaeological discourse, as well as the three digital reconstructions of the Maxentius mausoleum that we created. Furthermore, we provide an example software to demonstrate some of the possible concepts and techniques.
@@
Figure 1: This image illustrates the three proposed models of the Mausoleum of Maxentius discussed in this work. The first model (Maxentius A) closely follows Rasch’s reconstruction proposal [21]. The second model (Maxentius B) is based on Rasch’s model but features a closed dome and seven windows in the clerestory, modeled after the Tor de’ Schiavi [22]. The third model (Maxentius C) attempts to merge the dome and windows of Tor de’ Schiavi with the proportions of the Maxentius Mausoleum.
@@ II Background
The popularity of immersive technologies, such as AR and VR, for museums or expositions, teaching, and even research in archaeological contexts has grown [18, 19, 6, 23, 24]. As we focus on reconstructions in this work and its interdisciplinary nature, it’s important to clarify the different types of reconstructions in the digital realm. Digital reconstruction involves creating objects using computational tools, while virtual reconstruction is the use of virtual models to recover a building or manmade object visually [3]. Immersive reconstruction extends this further by embedding objects in environments along Milgram’s reality-virtuality continuum [5], such as AR or VR, to enhance user experience through interaction and context.
As is the nature of a reconstruction, it is possible that not all pieces of a monument, for instance, are known. Researchers must complement the known parts with modeled ones, whose form and structure derive from various sources, such as paintings, folklore, and derivations from existing pieces. These different sources need to be indicated, as Demetrescu describes in his work [9]. The sources of the reconstruction components, as well as the surrounding area and information on the object’s purpose or its material, are all data that can be transported via various means, especially so with the help of immersive technologies.
Immersive technologies can support reasoning [25] by embedding data into the physical world [26], the recall of data [27], or spatial judgment accuracy [28]. This is also relevant in virtual archaeology, as such environments have proven to create a heightened sense of presence, as already shown by Kyrlitsias, Christofi, Michael-Grigoriou, Banakou, and Ioannou [29]. Presence, as a form of psychological immersion, can also be called engagement [30], and is an important concept, as there is evidence that task performance and recall are boosted through heightened engagement [31].
The research direction that concerns itself with investigating how to exploit tools which engage the user to increase data understanding is called immersive analytics [11]. In their book on immersive analytics, Marriott et al. describe several additional relevant aspects apart from those already discussed. One of these is situated analytics, which places data representations in relation to relevant objects, places, and persons to enhance understanding and decision-making [32]. For example, it might overlay an existing object with analysis information about it and is thus tightly linked to AR applications. Another is the embodiment, which relates to the interaction of a human with a system through touch, gestures, or voice as a form of physical affordance, rather than the point-and-click interaction typical of desktop PCs [30]. That said, the concept of multi-modality is closely related, as it encompasses input and output addressing multiple senses, such as vision, hearing, touch, kinesthesis, and (less commonly) smell.
@@ III Maxentius Mausoleum
Figure 2: The current state of the Maxentius mausoleum. Only the crypt still remains. The front porch is partly missing and has partly been used to build a farmhouse.
@@
The monumental mausoleum was strongly connected to Maxentius’ rule and likely almost completed when Maxentius was killed in 312 AD [22, pp. 92–93]. Its few remaining traces provide an excellent case study for the application of virtual archaeology. Typologically, the mausoleum can be attributed to a type of podium round building, which – according to the architectural historian Jürgen Rasch – ultimately derives from the Pantheon. From 1977 to 1979, Rasch conducted an architectural survey of the site of the Maxentius Mausoleum and proposed a reconstruction of the building modeled on the Pantheon with its specific architectural components: a podium with a hexastyle front porch (pronaos), a domed circular building with seven niches, and a circular dome opening (oculus) for the second storey [34, p. 84]. However, this Pantheon version of the Maxentius Mausoleum has been questioned by Mark Johnson, referring to a temporally and typologically similar building – another late antique imperial mausoleum at the via Praenestina (the so-called Mausoleum of Tor de’ Schiavi) [34]. This monument is located, like the Maxentius Mausoleum, on the outskirts of Rome, close to an imperial villa suburbana. It follows the design of a podium round building featuring a slightly smaller, tetrastyle front porch with a domed vault without an oculus but four windows in the clerestory. These different perspectives, into which Maxentius’ Mausoleum can be classified, result in different conceivable models for its reconstruction, which will be discussed below.
@@ IV Immersive Analytics for the Maxentius Mausoleum
Regarding the Maxentius Mausoleum, we chose Rasch’s work as a foundation for a first reconstruction (Maxentius A), which closely follows his suggested measurements. The second model (Maxentius B) assumes a closed dome following Johnson’s proposal [22, pp. 91-92] and adds seven round windows in the clerestory with a diameter of 1.80 meters considering the proportions of Tor de’ Schiavi [34, p. 28]. The third model (Maxentius C) adapts the flatter dome of Tor de’ Schiavi [34, p. 68] and the windows with a larger diameter. This model visualizes, as suggested by Johnson, a version of the Maxentius Mausoleum modeled closely on Tor de’ Schiavi [22, p. 92]. Consequently, model B can be considered a merger of models A and C. Thus, we have three reconstruction models of the same object (see Fig. 1), which mainly differ in the design of the dome, allowing us to discuss which of the hypotheses about that part of the mausoleum is most plausible. We created the reconstructions with the 3D modeling software Blender. The components of the models consist of combined simple geometric forms, which were manipulated through manual node shifting to obtain necessary details. For each sub-component, once constructed from primitives, the measurements for those components were linearly scaled and applied. Thanks to the precise measurements obtained from Rasch’s work, the models provide correct proportions, which help us obtain a fitting impression of the mausoleum’s former glory. In Fig. 3, we demonstrate this via overlaying Maxentius A transparently over a photogrammetry-based model of the mausoleum acquired from airborne imaging using a drone.
Figure 3: Cross-faded version of Maxentius A. The different parts of the building are visible, and a photogrammetric recording of the current state of the crypt is embedded into the reconstruction.
We provide software (see Fig. 4) that presents the three hypothesized models of the Mausoleum of Maxentius. The software provides a VR environment, implemented with Unity C# and built to be used with an Oculus Quest 3 or Valve Index. We chose VR as it represents an s3D environment with head-tracking, which allows for improved spatial judgment. Inside the immersive environment, the user will start in front of the Maxentius A stairway. The user can move around, using the continuous movement metaphor, and can snap-turn to rotate.
@@
There is a panel with buttons to switch between the three reconstruction models and a toggle to turn on or off a coloring akin to the extended matrix by Demetrescu. Users can pick up this panel and carry it through the VR scene to change the representation at any position. Switching between models is especially relevant in a research context for comparing several different hypotheses directly while still in the iVR environment. Consequently, in dialogue with domain experts, compiling ways to compare the reconstruction models intuitively or conveniently helps to improve cognitive affordance.
To the right of the panel is a miniature of the model. By holding their virtual hands inside the miniature and performing a grabbing motion, a vibration of the controller indicates interaction, allowing them to control the angle of the light on the model. There is a small yellow sphere indicating the relative position of the light source to the miniature. This small sphere is intended to serve as a visualization to ease the intuition of where the light comes from.
This idiom is intended to aid in analyzing the source and distribution of sunlight for the mausoleum. For our mausoleum scenario, it would be beneficial to explore ways to effectively interact with light sources. An optimized light interaction system could lead to a more fluent experience, which in turn may help experts justify the different shapes of the dome and the light conditions inside. Furthermore, users can place small squares to leave small randomly colored markers at points of interest behind. In general, different types of markers could be developed. For example, users could be enabled to colorize specific areas of the wall to mark the locations of the numerous drawings on the walls, which were added long after the mausoleum was originally built. The type of marker, color coding, or even an acoustic signal could then be investigated for its effectiveness in locating the different spots.
Lastly, the software can also be used without a VR headset, utilizing very simple keyboard and mouse controls, allowing it to be viewed without the need for additional hardware. We provide the software via a Zenodo upload at [35].